CAVITY RING-DOWN LASER ABSORPTION SPECTROSCOPY OF THE E ^{3}Δ - X ^{3}Δ TRANSITION OF VN

\underline{TONGMEI MA}, J. W-H. LEUNG and A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

The (0,0) band of the electronic transition of VN at around 450.5 nm has been investigated using the technique of laser vaporization/reaction with free jet expansion and cavity ring-down laser absorption spectroscopy. A new transition system was observed, which has been designated as the E ^{3}Δ - X ^{3}Δ system. All three ΔΩ = 0 subband transitions were recorded and rotationally analyzed. Least squares fit of the measured line positions yielded molecular constants for the new E ^{3}Δ state. The bond length of the E^{3}Δ state was determined to be 1.6937 Å, which is the longest among the known states of VN. The E ^{3}Δ state is expected to arise from an electronic configuration 1δ^{1}10σ^{1} where the 10σ orbital is an antibonding orbital. A comparison of the observed electronic states of the isoelectronic TiO molecule supports the assignment.